# Elliptic Quantum Toroidal Algebra $U_{t_{1},t_{2},p}({\mathfrak{gl}}_{1,tor}),$ Vertex Operators and $L$-operators Hitoshi Konno and Andrey Smirnov

### v1

@@ 5.1 The type I case
${}^{(1,N-1)}{}_{-v/u}\langle 0|\Phi_{\lambda}(u){|0\rangle}_{v}^{(1,N)}{}$
$=\left(-\frac{(1-p)u}{v(\hbar^{1/2})^{N-1}}\right)^{n}\sum_{{\bf d}\in{\mathbb{N}}^{n}}(p^{-N}\mathfrak{z})^{-\sum_{a\in\lambda}d_{a}}\prod_{a\in\lambda}({\varphi^{\mu}_{\varsigma(a)}})^{N}\frac{(p^{-d_{a}}pu/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}{(p^{-d_{a}}\hbar u/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}{}$
×$\times\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}t_{1}^{-1}\frac{(p^{d_{a}-d_{b}}pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(p^{d_{a}-d_{b}}\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}{(p^{d_{a}-d_{b}}p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(p^{d_{a}-d_{b}}t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{\mu}.$ (5.3)
Let ${\cal N}_{\mu}$ be the the specialization of the integrand in (5.2) to $x_{a}=\varphi^{\mu}_{\varsigma(a)}\ (a\in\lambda)$ i.e. the ${\bf d}={\bf 0}$ term in (5.3). Then the $\mathrm{K}$-theoretic vertex function $V_{\lambda}^{\mu}(u,z)$ for $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$ is obtained as follows.
$V_{\lambda}^{\mu}(u,z)=\frac{1}{{\cal N}_{\mu}}{}^{(1,N-1)}{}_{-v/u}\langle 0|\Phi_{\lambda}(u){|0\rangle}_{v}^{(1,N)}$
=$\sum_{{\bf d}\in{\mathbb{N}}^{n}}(\hbar p^{-N-1}\mathfrak{z})^{-\sum_{a}d_{a}}\prod_{a\in\lambda}\frac{(\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}{(p\hbar^{-1}\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}\frac{(p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}{(pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}.$
@@
$V_{\lambda}^{\mu}(\hbar^{-1}u,\mathfrak{z}){}$
=$\sum_{{\bf d}\in{\mathbb{N}}^{n}}(\hbar p^{-N-1}\mathfrak{z})^{-\sum_{a}d_{a}}\prod_{a\in\lambda}\frac{(\hbar\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}{(p\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}\frac{(p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}{(pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}.$
@@
In the same way, the vertex function for ${\cal M}(n,r)$ is obtained from the vertex operator (4.24). We use the following formula obtained by combining Proposition 5.1 and the shuffle product formula (3.10).
$\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}\atop(a\in{\boldsymbol{\lambda}})}=\left(\hbar^{-(r-1)}\mathfrak{z}\right)^{-\sum_{a\in{\boldsymbol{\lambda}}}d_{a}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigr{|}_{{\boldsymbol{\mu}}},$
@@
$V_{\boldsymbol{\lambda}}^{{\boldsymbol{\mu}}}(u_{1}/\hbar,\cdots,u_{r}/\hbar,\mathfrak{z}):=\frac{1}{{{\cal N}}_{{\boldsymbol{\mu}}}}\ \quad{}^{(1,N-r)}{}_{(-\hbar)^{r}v/u_{1}\cdots u_{r}}\langle 0|\Phi_{\boldsymbol{\lambda}}(u_{1}/\hbar,\cdots,u_{r}/\hbar){|0\rangle}_{v}^{(1,N)}{}$
@@
×$\qquad\qquad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{(p\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}(t_{2}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}}{(pt_{1}^{-1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}(\hbar\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}}.$ (5.4)
@@
${\cal N}_{\boldsymbol{\mu}}=\prod_{1\leq i<j\leq r}<\Phi_{\emptyset}(u_{i})\Phi_{\emptyset}(u_{j})>\times\prod_{a\in{\boldsymbol{\lambda}}}\frac{(1-p)\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}}{v}\left(\frac{\hbar^{1/2}}{\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}}\right)^{-N}{}$
×$\quad\times\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\left(-\frac{\hbar^{-1/2}u_{i}}{\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}}\frac{(pu_{i}/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}{(\hbar u_{i}/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}\right)\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}t_{1}^{-1}\frac{(pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}{(p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}\ {}$
×$\qquad\qquad\qquad\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\Bigr{|}_{{\boldsymbol{\mu}}}.$
5.2 The type II dual case
@@
$\Psi^{*}_{\lambda^{(r)}}(u_{r})\cdots\Psi^{*}_{\lambda^{(1)}}(u_{1})$ $=$ $\prod_{1\leq i<j\leq r}<\Psi^{*}_{\emptyset}(u_{j})\Psi^{*}_{\emptyset}(u_{i})>\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\left(-\frac{\hbar^{1/2}u_{i}}{x_{a}}\frac{(\hbar p^{*}u_{i}/x_{a};p^{*})_{\infty}}{(u_{i}/x_{a};p^{*})_{\infty}}\right){}$
×$\ \times\prod_{i=1}^{r}\left((-u_{i})^{\alpha}e^{\Lambda_{0}^{\perp}-Q}\right)\prod_{a\in{\boldsymbol{\lambda}}}e^{h}\left(\frac{\hbar^{1/2}}{x_{a}}\right)^{c^{\perp}}:\prod_{a\in{\boldsymbol{\lambda}}}\widetilde{x}^{+}(x_{a})\prod_{i=1}^{r}\widetilde{\Psi}^{*}_{\emptyset}(u_{i}):,$
@@
[…]
@@
×$\qquad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{(p^{*}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}(t_{2}^{-1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}}{(p^{*}t_{1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}(\hbar^{-1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}}.$ (5.5)
@@
6 Exchange Relations of the Vertex Operators
We derive exchange relations of the vertex operators for $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$. We show that their coefficients are given by the elliptic dynamical instanton $R$-matrices, which arise as the transition matrices of the elliptic stable envelopes. These results can be easily extended to the case ${\cal M}(n,r)$.
6.1 Type I vertex operator
@@
$\Phi^{(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})\Phi^{(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})$ $=$ $\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right)\prod_{b\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(b)}}d_{p}x^{\prime\prime}_{b}\right):\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):\Phi_{\emptyset}(u_{1})\Phi_{\emptyset}(u_{2}){}$ (6.1)
×$\qquad\qquad\times\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}),$
where the chamber $\mathfrak{C}$ is given by $u_{1}\ll u_{2}$ with the stability condition $t_{1}/t_{2}>0$. Similarly we have the opposite order of composition
$\Phi^{(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})\Phi^{(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})$ $=$ $\prod_{b\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(b)}}d_{p}x^{\prime\prime}_{b}\right)\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right):\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):\Phi_{\emptyset}(u_{2})\Phi_{\emptyset}(u_{1}){}$ (6.2)
×$\qquad\qquad\times\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\overline{\mathfrak{C}},T^{1/2}}((\lambda^{\prime\prime},\lambda^{\prime});\mathfrak{z}),$
where the chamber $\overline{\mathfrak{C}}$ is given by $u_{1}\gg u_{2}$ with the stability condition $t_{1}/t_{2}>0$.
[…]
$R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}){|\beta^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\beta^{\prime\prime}\rangle}_{u_{2}}=\sum_{{\boldsymbol{\alpha}}\in({\cal P}^{2})_{n}}R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{{\boldsymbol{\beta}}}_{{\boldsymbol{\alpha}}}{|\alpha^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\alpha^{\prime\prime}\rangle}_{u_{2}},$ (6.3)
@@
$\widehat{\mathrm{Stab}}_{\overline{\mathfrak{C}},T^{1/2}}(\bar{{\boldsymbol{\alpha}}};\mathfrak{z})\bigr{|}_{\bar{{\boldsymbol{\gamma}}}}=\sum_{{\boldsymbol{\beta}}\in({\cal P}^{2})_{n}}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({{\boldsymbol{\beta}}};\mathfrak{z})\bigr{|}_{{{\boldsymbol{\gamma}}}}\ \bar{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})_{\boldsymbol{\alpha}}^{\boldsymbol{\beta}}.$ (6.5)
@@
By definition, the $R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})$ preserves the weight measured by $c^{\perp}$, i.e. the total number of the framing dimensions :
$[R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}),c^{\perp(1)}+c^{\perp(2)}]=0.$ (6.7)

### v2

@@ 5.1. The type I case
$\hskip-28.45274pt{}^{(1,N-1)}\hskip-19.91684pt{}_{-v/u}\langle 0|\Phi_{\lambda}(u){|0\rangle}_{v}^{(1,N)}{}$
$\hskip-28.45274pt=\left(-\frac{(1-p)u}{v(\hbar^{1/2})^{N-1}}\right)^{n}\sum_{{\bf d}\in{\mathbb{N}}^{n}}(p^{-N}\mathfrak{z})^{-\sum_{a\in\lambda}d_{a}}\prod_{a\in\lambda}({\varphi^{\mu}_{\varsigma(a)}})^{N}\frac{(p^{-d_{a}}pu/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}{(p^{-d_{a}}\hbar u/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}{}$
$\times\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}t_{1}^{-1}\frac{(p^{d_{a}-d_{b}}pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(p^{d_{a}-d_{b}}\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}{(p^{d_{a}-d_{b}}p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(p^{d_{a}-d_{b}}t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}{}$
$\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}(\lambda;\mathfrak{z})\bigr{|}_{\mu}.{}$
Let ${\mathcal{N}}_{\mu}$ be the the specialization of the integrand in (5.2) to $x_{a}=\varphi^{\mu}_{\varsigma(a)}\ (a\in\lambda)$ i.e. the ${\bf d}={\bf 0}$ term in (5.3). Then the $\mathrm{K}$-theoretic vertex function $V_{\lambda}^{\mu}(u,z)$ for $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$ is obtained as follows.
$\hskip-42.67912ptV_{\lambda}^{\mu}(u,z)=\frac{1}{{\mathcal{N}}_{\mu}}{}^{(1,N-1)}\hskip-19.91684pt{}_{-v/u}\langle 0|\Phi_{\lambda}(u){|0\rangle}_{v}^{(1,N)}$
$=\sum_{{\bf d}\in{\mathbb{N}}^{n}}(\hbar p^{-N-1}\mathfrak{z})^{-\sum_{a}d_{a}}\prod_{a\in\lambda}\frac{(\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}{(p\hbar^{-1}\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}{}$
$\quad\times\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}\frac{(p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}{(pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}.$
@@
$V_{\lambda}^{\mu}(\hbar^{-1}u,\mathfrak{z})=\sum_{{\bf d}\in{\mathbb{N}}^{n}}(\hbar p^{-N-1}\mathfrak{z})^{-\sum_{a}d_{a}}\prod_{a\in\lambda}\frac{(\hbar\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}{(p\varphi^{\mu}_{\varsigma(a)}/u;p)_{d_{a}}}{}$
$\qquad\qquad\times\prod_{a,b\in\lambda\atop\rho_{a}\not=\rho_{b}}\frac{(p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}{(pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}(\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{d_{a}-d_{b}}}.$
@@
In the same way, the vertex function for ${\mathcal{M}}(n,r)$ is obtained from the vertex operator (4.20). We use the following formula obtained by combining Proposition 5.1 and the shuffle product formula (3.9).
$\hskip-28.45274pt\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigr{|}_{x_{a}=\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}p^{d_{a}}\atop(a\in{\boldsymbol{\lambda}})}=\left(\hbar^{-(r-1)}\mathfrak{z}\right)^{-\sum_{a\in{\boldsymbol{\lambda}}}d_{a}}\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\bigr{|}_{{\boldsymbol{\mu}}},$
@@
$\hskip 5.69054ptV_{\boldsymbol{\lambda}}^{{\boldsymbol{\mu}}}(u_{1}/\hbar,\cdots,u_{r}/\hbar,\mathfrak{z})$ (5.4)
$:=\frac{1}{{{\mathcal{N}}}_{{\boldsymbol{\mu}}}}\ \quad{}^{(1,N-r)}\hskip-48.36958pt{}_{(-\hbar)^{r}v/u_{1}\cdots u_{r}}\langle 0|\Phi_{\boldsymbol{\lambda}}(u_{1}/\hbar,\cdots,u_{r}/\hbar){|0\rangle}_{v}^{(1,N)}{}$
@@
$\quad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{(p\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}(t_{2}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}}{(pt_{1}^{-1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}(\hbar\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p)_{d_{a}-d_{b}}}.{}$
@@
${\mathcal{N}}_{\boldsymbol{\mu}}=\prod_{1\leq i<j\leq r}<\Phi_{\emptyset}(u_{i})\Phi_{\emptyset}(u_{j})>\times\prod_{a\in{\boldsymbol{\lambda}}}\frac{(1-p)\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}}{v}\left(\frac{\hbar^{1/2}}{\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}}\right)^{-N}{}$
$\qquad\quad\times\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\left(-\frac{\hbar^{-1/2}u_{i}}{\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}}\frac{(pu_{i}/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}{(\hbar u_{i}/\varphi^{\mu}_{\varsigma(a)};p)_{\infty}}\right){}$
$\qquad\quad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}t_{1}^{-1}\frac{(pt_{1}^{-1}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(\hbar\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}{(p\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}(t_{2}\varphi^{\mu}_{\varsigma(a)}/\varphi^{\mu}_{\varsigma(b)};p)_{\infty}}\ {}$
$\qquad\qquad\qquad\times\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({\boldsymbol{\lambda}};\mathfrak{z})\Bigr{|}_{{\boldsymbol{\mu}}}.$
5.2. The type II dual case
@@
$\Psi^{*}_{\lambda^{(r)}}(u_{r})\cdots\Psi^{*}_{\lambda^{(1)}}(u_{1}){}$
$=\prod_{1\leq i<j\leq r}<\Psi^{*}_{\emptyset}(u_{j})\Psi^{*}_{\emptyset}(u_{i})>\prod_{i=1}^{r}\prod_{a\in{\boldsymbol{\lambda}}}\left(-\frac{\hbar^{1/2}u_{i}}{x_{a}}\frac{(\hbar p^{*}u_{i}/x_{a};p^{*})_{\infty}}{(u_{i}/x_{a};p^{*})_{\infty}}\right){}$
$\qquad\times:\prod_{a\in{\boldsymbol{\lambda}}}x^{+}(x_{a})\prod_{i=1}^{r}\Psi^{*}_{\emptyset}(u_{i}):,$
@@
$\hskip 8.53581ptV_{\boldsymbol{\lambda}}^{*{\boldsymbol{\mu}}}(u_{1},\cdots,u_{r},\mathfrak{z}^{*-1})$ (5.5)
$:=\frac{1}{{\mathcal{N}}^{*}_{{\boldsymbol{\mu}}}}\hskip 11.38092pt{}^{(1,N+r)}\hskip-42.67912pt{}_{(-)^{r}vu_{1}\cdots u_{r}}\langle 0|\Psi^{*}_{{\boldsymbol{\lambda}}}(u_{1},\cdots,u_{r}){|0\rangle}^{(1,N)}_{v}{}$
@@
$\quad\times\prod_{a,b\in{\boldsymbol{\lambda}}\atop\rho_{a}\not=\rho_{b}}\frac{(p^{*}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}(t_{2}^{-1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}}{(p^{*}t_{1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}(\hbar^{-1}\varphi^{\boldsymbol{\mu}}_{\varsigma(a)}/\varphi^{\boldsymbol{\mu}}_{\varsigma(b)};p^{*})_{d_{a}-d_{b}}}.{}$
@@
6. Exchange Relations of the Vertex Operators
We derive exchange relations of the vertex operators for $\mathrm{Hilb}^{n}({\mathbb{C}}^{2})$. We show that their coefficients are given by the elliptic dynamical instanton $R$-matrices, which arise as the transition matrices of the elliptic stable envelopes. These results can be easily extended to the case ${\mathcal{M}}(n,r)$.
6.1. The type I vertex operator
@@
$\hskip 11.38109pt\Phi^{(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})\Phi^{(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})$ (6.1)
$\hskip-14.22636pt=\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right)\prod_{b\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(b)}}d_{p}x^{\prime\prime}_{b}\right):\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):{}$
$\hskip-14.22636pt\times\Phi_{\emptyset}(u_{1})\Phi_{\emptyset}(u_{2})\hskip-8.53581pt\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}\hskip-11.38109pt<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}((\lambda^{\prime},\lambda^{\prime\prime});\mathfrak{z}),{}$
where the chamber $\mathfrak{C}$ is given by $|u_{1}|>|u_{2}|$ with the stability condition $t_{1}/t_{2}>0$. Similarly we have the opposite order of composition
$\hskip 8.53581pt\Phi^{(\mu^{\prime\prime})}_{\lambda^{\prime\prime}}(u_{2})\Phi^{(\mu^{\prime})}_{\lambda^{\prime}}(u_{1})$ (6.2)
$\hskip-14.22636pt=\prod_{b\in\lambda^{\prime\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime\prime}}_{\varsigma(b)}}d_{p}x^{\prime\prime}_{b}\right)\prod_{a\in\lambda^{\prime}}\left(\int_{0}^{\varphi^{\mu^{\prime}}_{\varsigma(a)}}d_{p}x^{\prime}_{a}\right):\prod_{a\in(\lambda^{\prime},\lambda^{\prime\prime})}x^{-}(x_{a}):{}$
$\hskip-14.22636pt\times\Phi_{\emptyset}(u_{2})\Phi_{\emptyset}(u_{1})\hskip-8.53581pt\prod_{a,b\in(\lambda^{\prime},\lambda^{\prime\prime})\atop\rho_{a}<\rho_{b}}\hskip-11.38109pt<x^{-}(x_{a})x^{-}(x_{b})>^{sym}\ \widehat{\mathrm{Stab}}_{\overline{\mathfrak{C}},T^{1/2}}((\lambda^{\prime\prime},\lambda^{\prime});\mathfrak{z}),{}$
where the chamber $\overline{\mathfrak{C}}$ is given by $|u_{1}|<|u_{2}|$ with the stability condition $t_{1}/t_{2}>0$.
Let $({\mathcal{P}}^{2})_{n}=\{{\boldsymbol{\alpha}}=(\alpha^{\prime},\alpha^{\prime\prime})\in{\mathcal{P}}\times{\mathcal{P}}\ |\ |{\boldsymbol{\alpha}}|=|\alpha^{\prime}|+|\alpha^{\prime\prime}|=n\ \}$. For ${\boldsymbol{\alpha}}=(\alpha^{\prime},\alpha^{\prime\prime})$, let $\bar{{\boldsymbol{\alpha}}}=(\alpha^{\prime\prime},\alpha^{\prime})$. We define the elliptic dynamical instanton $R$-matrix $R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})\in{\rm End}_{\mathbb{F}}({\mathcal{F}}^{(0,-1)}_{u_{1}}\widetilde{\otimes}{\mathcal{F}}^{(0,-1)}_{u_{2}})$ as the following transition matrix of the elliptic stable envelopes. For ${\boldsymbol{\alpha}},{\boldsymbol{\beta}},{\boldsymbol{\gamma}}\in({\mathcal{P}}^{2})_{n}$,
$R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}){|\beta^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\beta^{\prime\prime}\rangle}_{u_{2}}$ (6.3)
$=\sum_{{\boldsymbol{\alpha}}\in({\mathcal{P}}^{2})_{n}}R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})^{{\boldsymbol{\beta}}}_{{\boldsymbol{\alpha}}}{|\alpha^{\prime}\rangle}_{u_{1}}\widetilde{\otimes}{|\alpha^{\prime\prime}\rangle}_{u_{2}},{}$
@@
$\widehat{\mathrm{Stab}}_{\overline{\mathfrak{C}},T^{1/2}}(\bar{{\boldsymbol{\alpha}}};\mathfrak{z})\bigr{|}_{\bar{{\boldsymbol{\gamma}}}}$ (6.5)
$=\sum_{{\boldsymbol{\beta}}\in({\mathcal{P}}^{2})_{n}}\widehat{\mathrm{Stab}}_{\mathfrak{C},T^{1/2}}({{\boldsymbol{\beta}}};\mathfrak{z})\bigr{|}_{{{\boldsymbol{\gamma}}}}\ \bar{R}_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})_{\boldsymbol{\alpha}}^{\boldsymbol{\beta}}.{}$
@@
By definition, the $R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z})$ preserves the weight measured by $c$, i.e. the total number of the framing dimensions :
$[R_{T^{1/2}}(u_{1},u_{2};\mathfrak{z}),c^{(1)}+c^{(2)}]=0.$ (6.7)